OpenStudy (james1107):

What is the solution set of 2x 2 - 3x = 5? {-1, -5/2} {-1, 5/2} {1, -5/2}

OpenStudy (james1107):

@Ashes.boom @AloneS @skullpatrol @Elsa213 @rvc @fluttershyK @WhatTehHeck

OpenStudy (james1107):

@Zelda

OpenStudy (ashes.boom):

In mathematics, a solution set is the set of values that satisfy a given set of equations or inequalities. so the sets are (x, y) which means you can input the x values and see which ones work

alones (alones):

To me it looks like 2x^2−3x−5=5−5 2x^2−3x−5=0 Step 2: Factor left side of equation. (2x−5)(x+1)=0 Step 3: Set factors equal to 0. 2x−5=0 or x+1=0 x=5/2 or x=−1

OpenStudy (fluttershyk):

what answers do you know are wrong

OpenStudy (dayakar):

2x^2-3x-5=0 2x^2 +2x -5x -5 =0 [ split the middle term] 2x*x +2x*1 -5x -5*1=0 2x(x+1) -5(x+1) =0 (x+1) (2x_1) =0 solve it

OpenStudy (james1107):

So @AloneS it would be the last one

OpenStudy (fluttershyk):

@james3421 yes it would be @AloneS is right

OpenStudy (james1107):

thank yall

OpenStudy (james1107):

@AloneS plz medal @fluttershyK and i will medal you

alones (alones):

I already did xP

OpenStudy (james1107):

ok i medaled you

alones (alones):

O.M.G THANK YOU <3

OpenStudy (fluttershyk):

thanks guys yr awesome